# Topological Kleene Field Theories: A new model of computation

### v1

@@ Abstract.
In this article, we establish the foundations of a computational field theory, which we term Topological Kleene Field Theory (TKFT), inspired by Stephen Kleene’s seminal work on partial recursive functions. Our central result shows that any computable function can be simulated by the flow on a smooth bordism of a vector field with good local properties. More precisely, we prove that reaching functions on clean dynamical bordisms are exactly equivalent to computable functions, setting an alternative model of computation to Turing machines. The use of non-trivial topologies for the bordisms involved is essential for this equivalence, suggesting interesting connections between the topological structure of these flows and the computational complexity inherent in the functions. We emphasize that TKFT has the potential to surpass the computational complexity of both Turing machines and quantum computation.
@@
Primary: 37B10, Secondary: 03D10, 68Q04, 18F99
Ángel González-Prieto is supported by the Madrid Government (Comunidad de Madrid – Spain) under the Multiannual Agreement with the Universidad Complutense de Madrid in the line Research Incentive for Young PhDs, in the context of the V PRICIT (Regional Programme of Research and Technological Innovation) through the project PR27/21-029 and by the Ministerio de Ciencia e Innovación Project PID2021-124440NB-I00 (Spain). All authors are funded by the project “Computational, dynamical and geometrical complexity in fluid dynamics”, Ayudas Fundación BBVA a Proyectos de Investigación Científica 2021.
Eva Miranda is funded by the Catalan Institution for Research and Advanced Studies via an ICREA Academia Prize 2021 and by the Alexander Von Humboldt Foundation via a Friedrich Wilhelm Bessel Research Award and by the AGAUR project 2021 SGR 00603. She is also supported by the Spanish State Research Agency, through the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (project CEX2020-001084-M), and under grants reference PID2019-103849GB-I00 of AEI / 10.13039/501100011033 and PID2023-146936NB-I00 funded by MICIU/AEI/ 10.13039/501100011033 and, by ERDF/EU.
Daniel Peralta-Salas is supported by the grants CEX2023-001347-S, RED2022-134301-T and PID2022-136795NB-I0 funded by MCIN/AEI/ 10.13039/501100011033.
Significance statement
This work introduces Topological Kleene Field Theories (TKFTs), a novel framework that reinterprets the landscape of computation by means of flows on smooth bordisms. By bridging geometry and computation, TKFTs allow computable functions to be understood in terms of dynamical flows and topological decompositions. TKFT has the potential to go beyond the computational complexity of both Turing machines and quantum computing, opening new vistas in the theory of computation.
@@ 1. Introduction
One of the most interesting proposals in the model of partial recursive functions, as introduced by Kleene in [22], is that computability must be understood as output functions for Turing machines. Recall that, roughly speaking, a (binary) Turing machine $M$ is a device that receives as input a memory tape $t$ with finitely many symbols written on it, and uses a reading/writing head. At each step, the head reads the symbol on the tape $t$ beneath it and makes the decision to substitute it with another symbol and shift the head either to the right or to the left of the tape. The process continues until the machine $M$ reaches a halting state, or it continues indefinitely without stopping.
In this manner, given a natural number $n\in\mathbb{N}$, we can encode $n$ in the tape, for instance using its binary expansion. Then, we run the Turing machine $M$ with this tape. If $M$ halts at some point, the resulting tape can be understood as a natural number $f_{M}(n)\in\mathbb{N}$ encoded in binary. This output can be thus seen as a partial function $f_{M}:\mathbb{N}\dashrightarrow\mathbb{N}$, which is undefined on those numbers for which $M$ does not halt. Kleene proved that the set of partial functions $f_{M}$ arising from Turing machines can be described as composition of some basic and explicit functions, leading to a subset of partial functions called the set of partial recursive functions. This set is precisely the set of “computable functions” and constitutes an equivalent model of computation.
Therefore, given a partial recursive function $f:\mathbb{N}\dashrightarrow\mathbb{N}$, it is interesting to seek alternative ways of computing $f$. A natural approach to this problem is to consider dynamical systems induced by smooth vector fields on manifolds. Specifically, suppose that $W$ is a smooth bordism between two disjoint copies of a $2$-dimensional disc, denoted $\mathbb{D}_{1}$ and $\mathbb{D}_{2}$, i.e., $W$ is a smooth $3$-dimensional manifold with corners with boundary (the boundary contains $\mathbb{D}_{1}\sqcup\mathbb{D}_{2}$). We equip $W$ with a vector field $X$ transverse to the discs and tangent to the rest of the boundary, as in Figure 1.
@@
On each boundary disc $\mathbb{D}_{i}$, we can encode a copy of the natural numbers through an embedding $\mathbb{N}\hookrightarrow\mathbb{D}_{i}$, for instance, by encoding every natural number according to its binary expression, as in a square Cantor set. This defines a partial function $Z_{0}(W,X):\mathbb{N}\dashrightarrow\mathbb{N}$ by means of the ‘reaching function’ as follows: Given $n\in\mathbb{N}\subseteq\mathbb{D}_{1}$, we consider the flow of $X$ starting at $n$; if the flow hits $\mathbb{D}_{2}$ at a point $n^{\prime}\in\mathbb{N}\subseteq\mathbb{D}_{2}$, we set $Z_{0}(W,X)(n)=n^{\prime}$, otherwise the result is undefined. In this setting, the main result proven in this work is the following.
Theorem 1.1.
For any partial recursive function $f:\mathbb{N}\dashrightarrow\mathbb{N}$, there exists a clean dynamical bordism $(W,X)$ between discs, such that $X$ is a volume-preserving vector field and $Z_{0}(W,X)=f$. Conversely, the reaching function of any clean dynamical bordism is a partial recursive function. Accordingly, the category of clean dynamical bordisms is equivalent to that of computable functions.
In this theorem, “dynamical bordism” means that the vector field $X$ is transverse to the boundary discs $\mathbb{D}_{1}$ and $\mathbb{D}_{2}$, so it points inwards at $\mathbb{D}_{1}$ and outwards at $\mathbb{D}_{2}$, and it is tangent to the rest of the boundary. The “cleanness” property is a technical condition that ensures that the vector field is “well-behaved” in the sense that its zero set consists of finitely many points, and there is a “skeleton” where the flow of $X$ is generated by simple diffeomorphisms of discs (see Section 3.3 for precise definitions).
This result has several implications in the theory of dynamical systems. For instance, it shows that some problems in dynamics, such as determining whether the flow of a point will reach another point or even if the flow will be trapped, are undecidable due to the halting problem. Additionally, Theorem 1.1 means that the class of volume-preserving vector fields is Turing complete in the sense that any computable function can be simulated by the flow of a volume-preserving field on a bordism as above.
Even more interestingly, our construction relates the topology of the bordism to the complexity of the function it represents. As will be evident from our proof, a pair-of-pants as a bordism between one disc and two discs, will be understood as an ‘if’ in the computation, whereas a nontrivial handle will represent a ‘loop’ construction. In this sense, it is worth mentioning that this construction does require a non-trivial topology in the bordism to capture any partial recursive function. If we restrict ourselves to the class of bordisms obtained by choosing a diffeomorphism $\varphi:\mathbb{D}\to\mathbb{D}$ and considering the mapping cylinder $(W_{\varphi},X_{\varphi})$ induced by any homotopy between $\varphi$ and the identity, the reaching map is exactly $Z_{0}(W_{\varphi},X_{\varphi})=\varphi|_{\mathbb{N}}$ and thus it is continuous in the subspace topology for $\mathbb{N}\subseteq\mathbb{D}$; but not all the computable functions are continuous in this topology, as exemplified by the characteristic function of the even numbers. The reaching map of a general bordism may instead be discontinuous, providing the needed flexibility.
@@
The idea of investigating the interplay between dynamics and computation has been widely explored in the literature since the pioneering works of Moore [26, 27]. In this direction, dynamical systems have been shown to exhibit computational abilities, such as solving linear programming problems [9], undecidability in classical mechanics [17], the non-computability of Julia sets [8], the simulation of $\lambda$-calculus using a chemical-inspired machine [6], or the encoding of Turing machines in reaction-diffusion models [2]. This latter approach of embedding Turing machine dynamics into particular dynamical systems has been proven extremely useful in understanding the complexity of the solutions to differential equations, as in the context of analytic or polynomial families of ODEs [23, 24, 19, 7, 20, 21], in [34] for a potential well, or [35, 36] for the Euler equations modeling the dynamics of an ideal fluid flow.
@@
Inspired by these ideas, T. Tao suggested that the complexity of the Navier-Stokes equations modeling realistic fluids could be studied by understanding the encoding of universal Turing machines as part of the flow. This idea has been exploited further in [13, 11, 12, 14], where the authors proved that such universal Turing machines can be embedded within the flow of a steady state of the Euler equations in a $3$-dimensional manifold, leading to unexpected features in fluid dynamics. It is worth noticing that, despite the interest of these works, the approach taken in this paper is radically different: we do not aim to encode a Turing machine as a flow, which would require to be iterated many times to compute the final result; instead, we prove that the actual computable function can be simulated using a dynamical system in such a way that the outcome of the Turing machine can be represented in a single go of the flow. In other words, we do not encode the computability device as a flow, but the outcome function directly, so that the flow itself performs the whole computation process.
@@
In order to prove Theorem 1.1, we exploit an alternative representation of a Turing machine as a finite state machine. In this way, given a Turing machine $M_{f}$ computing a partial recursive function $f:\mathbb{N}\dashrightarrow\mathbb{N}$, we express $M_{f}$ as a directed finite graph where its nodes correspond to the states of $M_{f}$ and the edges represent the simple operations performed by the machine at each step, namely, reading a symbol, replacing it with another symbol and shifting the head. Encoding the tape as a point in the disc through a Cantor set-like embedding, we prove that these simple operations can be implemented using an explicit diffeomorphism. Hence, from the graph representing $M_{f}$, we create a CW-complex $W_{f}$ by placing a disc for each node and a mapping cylinder for each edge, representing the operation performed by the edge, as depicted in Figure 2. Using the natural vector field produced by a homotopy between the identity and each of the edge diffeomorphisms, we endow the CW-complex with a vector field. Then, by smoothing $W_{f}$ into a homotopically equivalent smooth manifold and doing a suitable (volume-preserving) extension of the flow, we get the desired bordisms whose dynamics represent the computable function $f$. A detailed proof of this result is provided in Appendix A and Theorem 3.15.
@@
It is worth mentioning that Theorem 1.1 can be reinterpreted from a categorical perspective, providing a computational analogue of Topological Quantum Field Theories (TQFTs) [38, 1, 39]. A TQFT is a phenomenological modelization of path integrals coming from quantum field theories independent of the underlying metric, in the form of a monoidal symmetric functor $Z:\textup{Bord}_{n}\to\textup{Vect}$ from the category of $n$-dimensional bordisms to the category of vector spaces. This functor must be seen as attaching to each $(n-1)$-manifold $M$ a vector space $Z(M)$ (the Hilbert space provided by a quantization of a classical system modeled by $M$) and, to each $n$-dimensional bordism $W$ between manifolds $M_{1}$ and $M_{2}$, it associates a linear map $Z(W):Z(M_{1})\to Z(M_{2})$ (the path integral induced by the quantum field theory between the Hilbert spaces of the boundaries).
Instead, in our context, the assignment $Z_{0}$ can be seen as a functor of $2$-categories $Z_{0}:\textup{Bord}_{3}^{\textup{dy}}\to\textup{PF}$ between the category of dynamical bordisms, i.e., bordisms with a well-behaved vector field, and the category of partial functions. In this language, Theorem 1.1 can be equivalently restated as follows.
Theorem 1.2.
On the subcategory $\textup{Bord}_{3}^{\textup{c-dy}}\subseteq\textup{Bord}_{3}^{\textup{dy}}$ of clean dynamical bordisms, the functor $Z_{0}$ restricts to a full and surjective on objects monoidal functor
$Z=Z_{0}|_{\textup{Bord}_{n}^{\textup{c-dy}}}:\textup{Bord}_{3}^{\textup{c-dy}}\longrightarrow\textup{PRF},$
onto the $2$-category of partial recursive functions.
@@
Due to its parallelism with Topological Quantum Field Theories, we shall refer to this functor $Z:\textup{Bord}_{n}^{\textup{c-dy}}\to\textup{PRF}$ as a Topological Kleene Field Theory (TKFT for short), named after Stephen Kleene who introduced partial recursive functions as a central piece in computation. Indeed, in his “Lectures on Computation”, Richard Feynman delves into the algebraic structures underpinning quantum computing, drawing parallels to Topological Quantum Field Theory (TQFT).
The TKFT represents a new model of computation in which computable functions are central objects that can be executed by following the flow of a vector field. Under this point of view, it opens new perspectives in complexity theory, since the computational complexity of the function can be related to the topological complexity of the bordism and the vector field. In fact, the bordisms used in Theorem 1.1 are forced to be non-trivial in the general case, precluding a deep relation between the topology and the complexity of the function to be represented.
@@
Furthermore, even though the proof developed in this work to establish Theorem 1.1 required the use of an auxiliary Turing machine, there may exist other clean dynamical bordisms with the same properties, but whose dynamics are constructed using different techniques. These non-Turing bordisms could provide a way of speeding-up the computation of certain partial recursive functions, outperforming the efficiency of classical computation. In this direction, it is then natural to ask whether TKFTs could offer more efficient approaches to computation than Turing machine-based computers or even if they could surpass non-deterministic methods, such as those developed in quantum computing.
Structure of the document
In Section 2, we review the fundamentals of partial recursive functions and Turing machines, with a special focus on the interpretation of Turing machines as finite state machines. In Section 3, we establish the framework of Topological Kleene Field Theories, as dynamical bordisms and the reaching functor $Z_{0}$. Section 4 provides a categorical interpretation of this result through the language of $2$-categories and functors. Section 5 compiles some insights about the capabilities of the proposed model and outlines some future research lines. Finally, the complete proof of the main Theorem 1.1 is presented in Appendix A.
2. Preliminaries: The Church-Turing Thesis
@@ 2.1. Turing Machines
This function $\Delta_{M}$ represents the intuitive idea that $t$ is a two-sided tape with finitely many non-black symbols and $M$ has a reader-writer head placed on the 00-th position of the tape. At each step, the head of $M$ reads the symbol under it and, according to the internal state $q$ of the machine, replaces it with the new symbol $s$ on the tape, shifts the head to the left ($\epsilon=-1$, or equivalently, the sequence to the right) or right ($\epsilon=1$, or the sequence to the left) and the machine changes its internal state to $q^{\prime}$. In this vein, given an initial tape $t$, the function $\Delta_{M}$ defines a dynamical system on $\mathcal{S}$ starting in $(t,q_{0})\in\mathcal{S}$. If, at some point, the system reaches a state of the form $(t^{\prime},q^{\prime})$ with $q^{\prime}\in Q_{\textrm{halt}}$, the dynamical system stops and we say that $M$ halts with tape $t^{\prime}$. Otherwise, we say that $M$ does not halt.

### v2

# Topological Kleene Field Theories as a model of computation

@@ Abstract.
In this article, we establish the foundations of a computational field theory, which we term Topological Kleene Field Theory (TKFT), inspired by Stephen Kleene’s seminal work on partial recursive functions and drawing parallels with Topological Field Theory. Our central result shows that any computable function can be simulated by the flow on a smooth bordism of a vector field with good local properties, setting an alternative model of computation to Turing machines. We thus establish that a computable function can be fully realized within a single go of a dynamical system, differing from previous works where computation is encoded as an iterative process. The output of the computable function emerges directly, laying the groundwork for potential applications that accelerate the physical realization of computation.
@@
Primary: 37B10, Secondary: 37C10, 03D10, 68Q04, 18N10
Corresponding author: Eva Miranda. Email: eva.miranda@upc.edu
@@ 1. Introduction
One of the most interesting proposals in the model of partial recursive functions, as introduced by Kleene in [23], is that computability must be understood as output functions for Turing machines. These are the mathematical abstractions of computers, comprised of a reading/writing head that manipulates a tape that works as an auxiliary memory. In this manner, a computable function is a partial function $f:\mathbb{N}\dashrightarrow\mathbb{N}$ that can be calculated by executing a Turing machine. Kleene proved that these computable functions coincide with partial recursive functions, which can be easily described as composition of some basic and explicit functions.
In this work, we show that it is possible to represent any computable function as the flow of a smooth vector field. Specifically, given such $f:\mathbb{N}\dashrightarrow\mathbb{N}$, we construct a smooth bordism $W_{f}$ with corners between two $2$-dimensional disks and a vector field $X$ transverse to the disks and tangent to the rest of the boundary (see Figure 1) such that the reaching function of the flow of $X$ at the boundary coincides with $f$.
@@
To be precise, on each boundary disk $\mathbb{D}$, we encode a copy of the natural numbers through an embedding $\mathbb{N}\hookrightarrow\mathbb{D}$, for instance, by encoding every natural number according to its binary expression, as in a square Cantor set. This defines a partial function $Z_{0}(W,X):\mathbb{N}\dashrightarrow\mathbb{N}$, called the ‘reaching function’, as follows: Given $n\in\mathbb{N}$ in the input disk, we consider the flow of $X$ starting at $n$; if the flow hits $\mathbb{D}_{2}$ at a point $n^{\prime}\in\mathbb{N}$ of the output disk, we set $Z_{0}(W,X)(n)=n^{\prime}$, otherwise the result is undefined.
Related work
@@
The idea of investigating the interplay between dynamics and computation has been widely explored in the literature since the pioneering works of Moore [26, 27]. In this direction, dynamical systems have been shown to exhibit computational abilities, such as solving linear programming problems [9], undecidability in classical mechanics [17], the non-computability of Julia sets [8], the simulation of $\lambda$-calculus using a chemical-inspired machine [6], the construction of Turing-complete cellular automata [40, 42, 41], or the encoding of Turing machines in reaction-diffusion models [2]. This latter approach of embedding Turing machine dynamics into particular dynamical systems has been proven useful in understanding the complexity of the solutions to differential equations, as in the context of analytic or polynomial families of ODEs [24, 25, 20, 7, 21, 22, 19], in [34] for a potential well, or [35, 36] for the Euler equations modeling the dynamics of an ideal fluid flow.
Inspired by these ideas, T. Tao suggested that the complexity of the Navier-Stokes equations modeling realistic fluids could be studied by understanding the encoding of universal Turing machines as part of the flow. This idea has been exploited further in [13, 11, 12, 14], where it was proven that such universal Turing machines can be embedded within the flow of a steady state of the Euler equations in a $3$-dimensional manifold, or in [18] for steady states of the Navier–Stokes equations via cosymplectic geometry.
@@
However, despite the interest of these works, the approach taken in this paper is deeply different: we do not aim to encode a Turing machine as a flow, which would require to be iterated many times to compute the final result; instead, we prove that the actual computable function can be simulated using a dynamical system in such a way that the outcome of the Turing machine can be represented in a single go of the flow.
Main result
The main theorem established in this work is the following.
Theorem A.
For any partial recursive function $f:\mathbb{N}\dashrightarrow\mathbb{N}$, there exists a clean dynamical bordism $(W,X)$ between disks, such that $X$ is a volume-preserving vector field and $Z_{0}(W,X)=f$. Conversely, the reaching function of any clean dynamical bordism is a partial recursive function. Accordingly, the category of clean dynamical bordisms is equivalent to that of computable functions.
In this theorem, “dynamical bordism” means that the vector field $X$ points inwards at the input disk and outwards at the output disc, and it is tangent to the rest of the boundary. The “cleanness” property is a technical condition that ensures that the vector field is “well-behaved” in the sense that its zero set consists of finitely many points, and there is a “skeleton” where the flow of $X$ is generated by simple diffeomorphisms of disks (see Section 3.3 for precise definitions).
It is worth noticing that the construction of Theorem A does require a non-trivial topology in the bordism to capture any partial recursive function. If we restrict ourselves to the class of bordisms obtained by taking the suspension of a diffeomorphism $\varphi:\mathbb{D}\to\mathbb{D}$ into its mapping cylinder $(W_{\varphi},X_{\varphi})$, the reaching map is exactly $\varphi|_{\mathbb{N}}$ and thus it is continuous in the subspace topology for $\mathbb{N}\subseteq\mathbb{D}$. But not all the computable functions are continuous in this topology. The reaching map of a general bordism may instead be discontinuous, providing the needed flexibility.
Sketch of the proof
@@
For the proof of Theorem A, we deeply exploit an alternative representation of a Turing machine as a finite state machine. In this way, given a Turing machine $M_{f}$ computing a partial recursive function $f:\mathbb{N}\dashrightarrow\mathbb{N}$, we express $M_{f}$ as a directed finite graph where its nodes correspond to the states of $M_{f}$ and the edges represent the simple operations performed by the machine at each step. Encoding the tape as a point in the disk through a Cantor set-like embedding, we prove that these simple operations can be implemented using an explicit diffeomorphism.
Hence, from the graph representing $M_{f}$, we create a CW-complex $W_{f}$ by placing a disk for each node and a mapping cylinder for each edge, representing the operation performed by the edge, as depicted in Figure 2. Using the natural vector field produced by a homotopy between the identity and each of the edge diffeomorphisms, we endow the CW-complex with a vector field. Then, by smoothing $W_{f}$ into a homotopically equivalent smooth manifold and doing a suitable (volume-preserving) extension of the flow, we get the desired bordisms whose dynamics represent the computable function $f$. A detailed proof of this result is provided in Section 4 and Theorem 5.9.
@@
We underscore that the thickening construction used in the proof of Theorem A is fully explicit and amenable, enabling effective simulation; cf. [30]. In this direction, Figure 3 shows a dynamical bordism simulating a partial recursive function, and different trajectories corresponding to four different inputs, depicted in different colours. The planes inside the bordism serve as computational checkpoints, each corresponding to a state of the underlying Turing machine.
Figure 3. Simulation of a partial recursive function through a dynamical bordism, reproduced from [30].
Implications
Theorem A has several consequences in the theory of dynamical systems. For instance, it shows that some problems in dynamics, such as determining whether the flow of a point will reach another point or even if the flow will be trapped, are undecidable due to the halting problem. Additionally, Theorem A means that the class of volume-preserving vector fields is Turing complete in the sense that any computable function can be simulated by the flow of a volume-preserving field on a bordism as above.
@@
It is worth mentioning that Theorem A can be reinterpreted from a categorical perspective, providing a computational analogue of Topological Quantum Field Theories (TQFTs) [38, 1, 39], as we shall explore in Section 5. Recall that TQFTs are a mathematical formulation of quantum field theories independent of the underlying metric purely in terms of category theory by using the category of bordisms. In this setting, we show that the reaching function $Z_{0}$ leads to a functor of $2$-categories between the category of dynamical bordisms and the category of partial functions. Due to this parallelism with Topological Quantum Field Theories, we decided to refer to the construction of Theorem A as a Topological Kleene Field Theory (TKFT for short). Indeed, in his “Lectures on Computation”, Richard Feynman delves into the algebraic structures underpinning quantum computing, drawing parallels to TQFT.
@@
TKFTs represent a new model of computation in which computable functions are central objects that can be executed by following the flow of a vector field. Under this point of view, it opens new perspectives in complexity theory, since the computational complexity of the function can be related to the topological complexity of the bordism and the vector field. Furthermore, even though the proof developed in this work to establish Theorem A requires the use of an auxiliary Turing machine, there may exist other clean dynamical bordisms with the same properties, but whose dynamics are constructed using different techniques. These non-Turing bordisms could provide a way of speeding-up the computation of certain partial recursive functions, outperforming the efficiency of classical computation, with a view towards a physically realizable beyond-Turing model of calculation.
Acknowledgments
The authors were supported by the Spanish State Research Agency through projects PID2021-124440NB-I00, PID2023-146936NB-I00, RED2022-134301-T, and project PID2022-136795NB-I00 under MCIN/AEI/10.13039/501100011033 and ERDF/EU. EM was supported by the Catalan Institution for Research and Advanced Studies through an ICREA Academia Prize 2021. All authors were partially supported by the project Computational, Dynamical and Geometrical Complexity in Fluid Dynamics (COMPLEXFLUIDS), Fundación BBVA 2021, and the bilateral AEI–DFG projects PCI2024-155042-2 and PCI2024-155062-2. This work also benefited from the Severo Ochoa and María de Maeztu Program for Centres and Units of Excellence in R&D (CEX2020-001084-M and CEX2023-001347-S) of the Spanish State Research Agency.
2. The Church-Turing Thesis
@@ 2.1. Turing Machines
This function $\Delta_{M}$ represents the intuitive idea that $t$ is a two-sided tape with finitely many non-black symbols and $M$ has a reader-writer head placed on the 0-th position of the tape. At each step, the head of $M$ reads the symbol under it and, according to the internal state $q$ of the machine, replaces it with the new symbol $s$ on the tape, shifts the head to the left ($\epsilon=-1$, or equivalently, the sequence to the right) or right ($\epsilon=1$, or the sequence to the left) and the machine changes its internal state to $q^{\prime}$. In this vein, given an initial tape $t$, the function $\Delta_{M}$ defines a dynamical system on $\mathcal{S}$ starting in $(t,q_{0})\in\mathcal{S}$. If, at some point, the system reaches a state of the form $(t^{\prime},q^{\prime})$ with $q^{\prime}\in Q_{\textrm{halt}}$, the dynamical system stops and we say that $M$ halts with tape $t^{\prime}$. Otherwise, we say that $M$ does not halt.